-5x + 6 = -115 + 6x

Answer by checkley77(12844)

-5x + 6 = -115 + 6x
-5x-6x=-115-6
-11x=-121
x=121/-11
x=11 ans.
Proof:
-5*11=6=-115+6*11
-55+6=-115+66
-49=49